# Vision-Based Power Line Cables and Pylons Detection for Low Flying Aircrafts

### v1

@@ Abstract
Power lines are dangerous for low-flying aircrafts, especially in low-visibility conditions. Thus, a vision-based system able to analyze the aircraft’s surroundings and to provide the pilots with a “second pair of eyes” can contribute to enhancing their safety. To this end, we have developed a deep learning approach to jointly detect power line cables and pylons from images captured at distances of several hundred meters by aircraft-mounted cameras. In doing so, we have combined a modern convolutional architecture with transfer learning and a loss function adapted to curvilinear structure delineation. We use a single network for both detection tasks and demonstrated its performance on two benchmarking datasets. We have integrated it within an onboard system and run it in flight, and have demonstrated with our experiments that it outperforms the prior distant cable detection method [1] on both datasets, while also successfully detecting pylons, given their annotations are available for the data.
@@ 1 Introduction
Low-flying aircrafts, such as light planes and helicopters, are always at risk of colliding with ground obstacles. Power lines are especially dangerous because they are hard to see, even in good weather. In 2010, the Federal Aviation Administration used 13 years worth of accident data from the National Transportation Safety Board (NTSB) database to estimate that there are 76 collisions with power line per year on average and that 30% of them resulted in fatalities [4]. Utility poles and transmission towers, which we will collectively refer to as pylons, constitute additional dangers and are often found in the vicinity of power line cables. Many of the accidents stem from the fact that pilots are expected to simultaneously carry out many tasks—steering the aircraft, navigating, monitoring the weather, and performing mission-related tasks—that all require a high level of concentration. This sometimes causes cognitive overload and loss of situational awareness [5]. An automated system that would alert the pilots to an impending collision danger would therefore be extremely valuable.
@@
Systems, such as the FLARM [6], which relies on a database of known obstacles, are a good start. However, there always is the possibility of obstacles not being in the database. Thus, there has long been interest in developing video-based algorithms to detect power line cables. Most early work relied on image processing methods designed for line detection, both for detection of cables [7, 8, 9, 10] and pylons [11]. Feature-based machine learning techniques then started being used [8, 12, 13] but were superseded by deep-learning based methods, both for power line cables [14, 15, 16, 17, 18] and transmission towers [19]. Most of prior research focuses on either the detection of power line cables or that of pylons [20] and does not treat them jointly. Methods for power line cables detection are frequently developed in the context of power line infrastructure inspection, where the cables are likely to be more clearly visible than at larger distances encountered during regular small aircraft flights. In this work we tackle this latter challenging scenario with the aim of developing an aircraft safety system assisting the pilots in detecting these notorious dangers.
@@
We propose a new deep-learning based approach to jointly detect power line cables and pylons. We use a pre-trained ConvNext [21] as our backbone and formulate the detection task as one of regressing a map of distances to the obstacles from the images. To validate our method we compare with approach proposed by Stambler et al. [1] and use the accompanying NE-VBW dataset – a public high-resolution dataset created for vision-based power line cables detection in the context of developing aircraft safety system. We demonstrate that combining transfer learning, a modern backbone, and a suitable loss function designed for curvilinear structure detection boosts performance beyond the previously proposed method [1] on their dataset.
@@
There are other datasets that could be used to evaluate cable detection algorithms [10, 22, 23, 24, 25, 26, 27, 28]. However, they have various limitations, including: a viewing perspective from the level close to the ground and mostly in urban settings [10, 28], scaled down higher resolution images with only image-level labels [22, 23], small dataset size [27] or well visible power lines [25, 24], including scenarios intended for testing cable inspection methods with power lines seen from a short distance [26]. We are facing a very different scenario with broad fields of view, distant barely visible power lines and high-resolution images. To develop our detection method we created our more diverse DDLN dataset, recorded with short focals that provide information about a large area in front of the aircraft. We also evaluate our approach on it and demonstrate the strong performance of our approach. Finally, we show that, given additional annotations for pylons, a single network can easily and successfully be trained to detect them in addition to cables. Furthermore, when only pylon detection is needed and small detection errors are tolerated, it performs competitively to the state-of-the-art real-time object detector [3] of similar size.
@@
In the following section 2 we present prior works on vision-based power line cables and pylons detection. In section 3 we describe our model architecture and the loss function. Section 4 presents our results, preceded by the descriptions of the datasets used for the evaluation and the implementation details. Finally, we summarize our work and discuss the limitations of the presented method together with the possible future directions in section 5.
@@ 2.1 Detecting Power Lines
Another line of research focused on using object, line or key points detectors. Nguyen et al. [36] proposed LS-Net, a single-shot detector, combining classification and line endpoints regression tasks, performed over a set of four overlapping grids of cells. Their method has been evaluated on a dataset [24] featuring 800 crops extracted from full HD frames, scaled to $512\times 512$ px patches, with power line cables clearly visible over the terrain background. Abdelfattah et al. [26] introduced the TTPLA dataset, featuring high-resolution aerial images captured by a drone flying within a close distance from power line infrustructure and annotated with cable and transmission tower masks. They also established an instance segmentation baseline with the use of YOLACT [37], an extension of single-shot object detector with a segmentation capability. Recently, Xing et al. [38] designed a power line detection method for use by inspection drones. They used YOLO [39], trained solely on synthetic data which additionally classified the inclination of a power line cable within the detected bounding box. However, as in [26], their data was restricted to close views of power line cables from the perspective of a drone hovering above the inspected infrustructure. Finally, a key point detection scheme was proposed by Dai et al. [40]. They formulate the detection as a regression of 5 key points evenly distributed across each cable in the input image. The points regressed by the network are subsequently clustered and then a curve representing the complete power line cable can be fitted to their locations. Their method was evaluated on PLDU and PLDM datasets [25], which contain in total 860 images with $540\times 360$ px resolution, each of which with several wires to be detected. These datasets feature again very (PLDU) and relatively (PLDM) close views of cables, often clearly visible against an urban (PLDU) or a mountainous (PLDM) scenery. This makes the cable detection task easier than in our challenging setting of detections at significant distances and over complete high-resolution image.
@@
Other more recent segmentation approaches to power line cables detection included models based on UNet [44] or UNeXt [45] architectures. They leverage generative adversarial networks (GANs) trained to highlight power lines in the images [46] or added transformer blocks to CNN feature extractors enabling global attention [47]. Most of these methods [45, 46, 47] were evaluated on the TTPLA dataset [26], which features high-resolution close views of cables and pylons, mostly suitable for power line infrastructure inspection. Jaffari et al. [44] evaluated their model on IVRL [24] and PLDU [25], both containing well visible cables crossing the images and the former with low variability of wire appearance. Cheng et al. [45] additionally measured their performance on the VITL dataset [27] which is rather scarce with only 400 crops of higher-resolution images, with a size of $256\times 256$ pixels each. In this work, we evaluated the performance of the models on two datasets – NE-VBW [1] and DDLN dataset collected by us – both of which comprise complete, high-resolution images captured from the flying aircrafts within significant distances from the detected cables and pylons, a setting relevant in the development of a system aiming to improve the aircraft safety.
Cable segmentation models have also been used in the field of computational photography, along with a wire segmentation and in-painting model trained to improve the image aesthetics by removing cables from the photographs. Chiu et al. [28] have collected the WireSegHR dataset containing 6000 very high-resolution images with pixel-wise annotations of wire locations. Judging by the released test set of 420 images, the dataset contains various kinds of wires, cables along with other elongated objects, often with challenging appearance and poor visibility in a diverse set of urban sceneries and in cluttered scenes. The images were captured from the ground level as in a typical photography setting. This makes for a different viewing perspective than the one encountered by aircraft. In the images from our DDLN dataset and from the NE-VBW dataset [1], on which we evaluated our method, the wires are mostly seen by the aircraft-mounted camera during flight.
@@ 3 Method
For cable detection, we differ from the state-of-the-art approach of Stambler et al. [1] by predicting only the undirected distance to the closest wire, as opposed to distance, direction, and a binary value indicating the presence or absence of a cable. The superior performance of our model, which we demonstrate in the result section, is achieved by combining transfer learning, a modern convolutional backbone, and a loss function that favors cable-like topologies.
@@ 3.1 Network Architecture
Figure 1: Architecture of the joint power lines and poles detection model. Output heads regress cables and poles distance masks directly from the common feature map $F{}$ generated by the fine-tuned truncated ConvNeXt-Tiny feature extractor.
@@ 3.2 Loss Function
where $N$ is the total number of pixels in $\mathbf{x}^{k}$ and the $\mathbf{x}^{k}$ distance masks are quantized to integer values. We use $f$ to write the weighted version of MSD
@@ 4.1 Datasets
There are many publicly available annotated datasets that can be used to validate power line detection algorithms. They include USF [10], IVRL [22, 23, 24], PLDU and PLDM [25], TTPLA [26], VITL [27] and WireSegHR [28]. As previoussly discussed in section 1, these have their shortcomings, including: viewing perspective and setting different from the one encountered in manned aircraft flights [10, 28], small size of the dataset with crops of larger images [27], containing scaled down larger images, with low resolution and image-level labels [22, 23] or featuring clearly visible power lines [24, 25, 26], oftenly seen from a short distance. The latter are more suitable to test algorithms designed to guarantee the safety of UAVs operating in the proximity of the power line infrastructure, mostly for inspection purposes. However, in this work, we want to perform the detection at longer ranges for collision-avoidance purposes. Furthermore, we are doing it in images captured by aircraft-mounted cameras with short focal lengths to cover the large area in front of the aircraft required to ensure safety.
@@
NE-VBW Dataset [1]. It is a public dataset containing 109 images of images of size up to $6576\times 4384$ px. They were extracted from recordings captured by aircraft-mounted, forward facing cameras with a large focal length. The high resolution and the large focal length of the cameras allowed for the capture of power line cables at long ranges within a narrow field of view – 25° horizontally and 18° vertically. The images were labeled using pixel-wise masks of power line cables. All data originates from flights conducted around Pittsburgh, PA in the USA.
DDLN Dataset. We built it ourselves from videos acquired during flights over Switzerland, the UK, Brazil, and the USA. It contains 865 images of size $4096\times 3000$ px each. They were recorded by aircraft-mounted, forward facing cameras with a wide field of view – up to 81° horizontally and 64° vertically. They are significantly larger than those of the NE-VBW dataset. This makes the detection task more challenging but provides more context by covering a larger area in front of the aircraft. The videos were recorded at the rate of 6 frames per second, but the annotated frames were extracted with sufficient time intervals between them to provide for a diversified scene content. Polyline approximations of individual power line cables were generated by hand and bounding boxes were used to denote the pylons. Additionally, exclusion areas, that is, regions containing thin objects not associated with cables—flagpoles, cranes, communication towers—were marked by bounding boxes. We did not use them to create training data and excluded them from the pylon detection performance evaluation.
@@ 4.2 Baselines
PIDNet-S [2] – cable detection. This is a family of models designed for semantic segmentation. We train it without boundary-aware loss components, thus also without D branch supervision. We did this because these components are mostly useful for multi-class semantic segmentation with multiple, potentially confusing, object boundaries in the image. These issues do not arise for power line cables. We start from the ImageNet-pretrained checkpoint of the PIDNet-S model and add at the end of it a single $2\times 2$ strided convolution layer to match the downsampling factor of our models. We also replaced the SGD optimizer with Adam [83] and did not use the poly learning rate strategy.
RT-DETR [3] – pylon detection. This object detector uses the ResNet18 backbone [84]. As PIDNet-S, it has been designed with real-time applications in mind. It builds on top of transformer-based object detectors. We additionally combine it with SAHI [74], training and inference strategy for object detection in high-resolution images. We evaluate two variants of patched input scales. In both cases, we run the object detector over full-resolution image and its split into $1024\times 1024$ overlapping patches, appropriately resizing the inputs for the detector and combining all of its outputs. In the second variant we additionally use similarly generated set of $512\times 512$ and $256\times 256$ patches and add their detection results to the output. For evaluation, we rescale the final predicted bounding boxes to the downsampled output resolution of our model, use them to generate binary segmentation masks and convert them into distance masks. Also, for this model we initialize it with the pretrained weights provided with the official model implementation, obtained from training on the Objects365 [85] dataset and subsequent fine-tuning on the COCO [86] dataset.

### v3

@@ Abstract
Power lines are dangerous for low-flying aircrafts, especially in low-visibility conditions. Thus, a vision-based system able to analyze the aircraft’s surroundings and to provide the pilots with a “second pair of eyes” can contribute to enhancing their safety. To this end, we develop a deep learning approach to jointly detect power line cables and pylons from images captured at distances of several hundred meters by aircraft-mounted cameras. In doing so, we combine a modern convolutional architecture with transfer learning and a loss function adapted to curvilinear structure delineation. We use a single network for both detection tasks and demonstrate its performance on two benchmarking datasets. We have also integrated it within an onboard system and run it in flight. We show with our experiments that it outperforms the prior distant cable detection method [1] on both datasets, while also successfully detecting pylons, given their annotations are available for the data.
@@ 1 Introduction
Low-flying aircraft, such as light planes and helicopters, are always at risk of colliding with ground obstacles. Power lines are especially dangerous because they are hard to see, even in good weather. In 2010, the Federal Aviation Administration used 13 years worth of accident data from the National Transportation Safety Board (NTSB) database to estimate that there are 76 collisions with power line per year on average and that 30% of them resulted in fatalities [2]. Utility poles and transmission towers, which we will collectively refer to as pylons, constitute additional dangers and are often found in the vicinity of power line cables. Many of the accidents stem from the fact that pilots are expected to simultaneously carry out many tasks—steering the aircraft, navigating, monitoring the weather, and performing mission-related tasks—that all require a high level of concentration. This sometimes causes cognitive overload and loss of situational awareness [3]. An automated system that would alert the pilots to an impending collision danger would therefore be extremely valuable.
@@
Systems that rely on a database of known obstacles, such as FLARM [4], are a good start. However, there always is the possibility of obstacles not being in the database. Thus, there has long been interest in developing video-based algorithms to detect power line cables. Most early work relied on image processing methods designed for line detection, both for detection of cables [5, 6, 7, 8] and pylons [9]. Feature-based machine learning techniques then started being used [6, 10, 11] but were superseded by deep-learning based methods, both for power line cables [12, 13, 14, 15, 16] and transmission towers [17]. Most of prior research focuses on either the detection of power line cables or that of pylons [18] and does not treat them jointly. Methods for power line cables detection are frequently developed in the context of power line infrastructure inspection, where the cables are likely to be more clearly visible than at larger distances encountered during regular small aircraft flights. In this work we tackle this latter challenging scenario with the aim of developing an aircraft safety system assisting the pilots in detecting these notorious dangers.
@@
We propose a new deep-learning based approach to jointly detect power line cables and pylons. We use a pre-trained ConvNeXt [19] as our backbone and formulate the detection task as one of regressing a map of distances to the obstacles from the images. To validate our method we compare it with the approach proposed by Stambler et al. [1] and use the accompanying NE-VBW dataset – a public high-resolution dataset created for vision-based power line cables detection in the context of developing aircraft safety system. We demonstrate that combining transfer learning, a modern backbone, and a suitable loss function designed for curvilinear structure detection boosts performance beyond the previously proposed method [1] on their dataset.
@@
There are other datasets that could be used to evaluate cable detection algorithms [8, 20, 21, 22, 23, 24, 25, 26]. However, they have various limitations, including: a viewing perspective from the level close to the ground and mostly in urban settings [8, 26], scaled down higher resolution images with only image-level labels [20, 21], small dataset size [25] or well visible power lines [23, 22], including scenarios intended for testing cable inspection methods with power lines seen from a short distance [24]. We are facing a very different scenario with broad fields of view, distant barely visible power lines and high-resolution images. To develop our detection method we created our more diverse DDLN dataset, recorded with short focals that provide information about a large area in front of the aircraft. We also evaluate our approach on it and demonstrate the strong performance of the proposed method. Finally, we show that, given additional annotations for pylons, a single network can easily and successfully be trained to detect them in addition to cables. Furthermore, when only pylon detection is needed and small detection errors are tolerated, it performs competitively to the state-of-the-art real-time object detector [27] of similar size.
@@
In the following section 2 we present prior works on vision-based power line cables and pylons detection. In section 3 we describe our model architecture and the loss function. Section 4 presents our results, accompanied by the descriptions of the datasets used for the evaluation, baselines we compare with and the implementation details. Finally, we summarize our work and discuss the limitations of the presented method together with the possible future directions in section 5.
@@ 2.1 Detecting Power Lines
Another line of research focused on using object, line or key points detectors. Nguyen et al. [35] proposed LS-Net, a single-shot detector, combining classification and line endpoints regression tasks, performed over a set of four overlapping grids of cells. Their method has been evaluated on a dataset [22] featuring 800 crops extracted from full HD frames, scaled to $512\times 512$ px patches, with power line cables clearly visible over the terrain background. Abdelfattah et al. [24] introduced the TTPLA dataset, featuring high-resolution aerial images captured by a drone flying within a close distance from power line infrastructure and annotated with cable and transmission tower masks. They also established an instance segmentation baseline with the use of YOLACT [36], an extension of single-shot object detector with a segmentation capability. Recently, Xing et al. [37] designed a power line detection method for use by inspection drones. They used YOLO [38], trained solely on synthetic data which additionally classified the inclination of a power line cable within the detected bounding box. However, as in [24], their data was restricted to close views of power line cables from the perspective of a drone hovering above the inspected infrastructure. Finally, a key point detection scheme was proposed by Dai et al. [39]. They formulate the detection as a regression of 5 key points evenly distributed across each cable in the input image. The points regressed by the network are subsequently clustered and then a curve representing the complete power line cable can be fitted to their locations. Their method was evaluated on PLDU and PLDM datasets [23], which contain in total 860 images with $540\times 360$ px resolution, each of which with several wires to be detected. These datasets feature again very (PLDU) and relatively (PLDM) close views of cables, often clearly visible against an urban (PLDU) or a mountainous (PLDM) scenery. This makes the cable detection task easier than in our challenging setting of detections at significant distances and over complete high-resolution image.
@@
Other more recent segmentation approaches to power line cables detection included models based on UNet [43] or UNeXt [44] architectures. Others leverage generative adversarial networks (GANs) trained to highlight power lines in the images [45] or added transformer blocks to CNN feature extractors enabling global attention [46]. Most of these methods [44, 45, 46] were evaluated on the TTPLA dataset [24], which features high-resolution close views of cables and pylons, mostly suitable for power line infrastructure inspection. Jaffari et al. [43] evaluated their model on IVRL [22] and PLDU [23], both containing well visible cables crossing the images and the former with low variability of wire appearance. Cheng et al. [44] additionally measured their performance on the VITL dataset [25] which is rather scarce with only 400 crops of higher-resolution images, with a size of $256\times 256$ pixels each. In this work, we evaluated the performance of the models on two datasets – NE-VBW [1] and DDLN dataset collected by us – both of which comprise complete, high-resolution images captured from the flying aircraft within significant distances from the detected cables and pylons, a setting relevant in the development of a system aiming to improve the aircraft safety.
Cable segmentation models have also been used in the field of computational photography, along with a wire segmentation and in-painting model trained to improve the image aesthetics by removing cables from the photographs. Chiu et al. [26] have collected the WireSegHR dataset containing 6000 very high-resolution images with pixel-wise annotations of wire locations. Judging by the released test set of 420 images, the dataset contains various kinds of wires and cables, along with other elongated objects, often with challenging appearance and poor visibility in a diverse set of urban sceneries and in cluttered scenes. The images were captured from the ground level as in a typical photography setting. This makes for a different viewing perspective than the one encountered by aircraft. In the images from our DDLN dataset and from the NE-VBW dataset [1], on which we evaluated our method, the wires are mostly seen by the aircraft-mounted camera during flight.
@@ 3 Method
For cable detection, we differ from the approach of Stambler et al. [1] by predicting only the undirected distance to the closest wire, as opposed to distance, direction, and a binary value indicating the presence or absence of a cable. The superior performance of our model, which we demonstrate in the result section, is achieved by combining transfer learning, a modern convolutional backbone, and a loss function that favors cable-like topologies.
@@ 3.1 Network Architecture
Figure 1: Architecture of the joint power line cables and pylons detection model. Output heads regress cables and pylons distance masks directly from the common feature map $F{}$ generated by the fine-tuned truncated ConvNeXt-Tiny feature extractor.
@@ 3.2 Loss Function
where $N$ is the total number of pixels in $\mathbf{x}^{k}$ and the $\mathbf{x}^{k}$ distance masks are discretized to integer values. We use $f$ to write the weighted version of MSD
@@
3.3 System Deployment
We deployed DDLN-S, our smallest model variant for joint cables and pylons detection, to Eval Kit, Daedalean’s hardware flight assistant system and set it up on a Robinson R44 Raven II helicopter. In order to achieve that, the PyTorch representation of the model and its trained weights were first converted into the ONNX format and then optimized using TensorRT API for the target system and an NVidia GeForce 3060 GPU card. To increase the inference speed, we additionally used half-precision floating point format for the network.
We also developed a pre- and post-processing pipelines, applied to the model’s inputs and to its predictions, respectively. The $4096\times 3000$ px raw images captured by the camera are debayered, padded and split into $1024\times 1024$ px non-overlapping patches. The patches are processed by the GPU in batches of size 4 and then stitched into a $128\times 93$ px output representing detections over the complete camera frame. In parallel, dense inverse search optical flow [82, 83] is applied to the two subsequent downsampled camera frames to estimate the movement in the image. Once both results are available, the previous model prediction for the complete frame is warped according to the optical flow output and then it is combined with the prediction for the current frame by computing the average of the two. This averaging operation aims to make the result less susceptible to both false positive and false negative noise. In the following step, we threshold the averaged distance masks turning cable and pylon detections into segmentation masks. Finally, these binary masks are upscaled to the original image size, overlapped with the input image and the result is forwarded to the system’s user interface.
@@ 4.1 Datasets
There are many publicly available annotated datasets that can be used to validate power line detection algorithms. They include USF [8], IVRL [20, 21, 22], PLDU and PLDM [23], TTPLA [24], VITL [25] and WireSegHR [26]. As previously discussed in section 1, these have their shortcomings, including: viewing perspective and setting different from the one encountered in manned aircraft flights [8, 26], small size of the dataset with crops of larger images [25], containing scaled down larger images, with low resolution and image-level labels [20, 21] or featuring clearly visible power lines [22, 23, 24], oftenly seen from a short distance. The latter are more suitable to test algorithms designed to guarantee the safety of UAVs operating in the proximity of the power line infrastructure, mostly for inspection purposes. However, in this work, we want to perform the detection at longer ranges for collision-avoidance purposes. Furthermore, we are doing it in images captured by aircraft-mounted cameras with short focal lengths to cover the large area in front of the aircraft required to ensure safety.
@@
NE-VBW Dataset [1]. It is a public dataset containing 109 images of size up to $6576\times 4384$ px. They were extracted from recordings captured by aircraft-mounted, forward facing cameras with a large focal length. The high resolution and the large focal length of the cameras allowed for the capture of power line cables at long ranges within a narrow field of view – 25° horizontally and 18° vertically. The images were labeled using pixel-wise masks of power line cables. All data originates from flights conducted around Pittsburgh, PA in the USA.
DDLN Dataset. We built it ourselves from videos acquired during flights over Switzerland, the UK, Brazil, and the USA. It contains 865 images of size $4096\times 3000$ px each. They were recorded by aircraft-mounted, forward facing cameras with a wide field of view – up to 81° horizontally and 64° vertically. It is significantly wider than the one in the images from the NE-VBW dataset. This makes the detection task more challenging but provides more context by covering a larger area in front of the aircraft. The videos were recorded at the rate of 6 frames per second, but the annotated frames were extracted with sufficient time intervals between them to provide for a diversified scene content. Polyline approximations of individual power line cables were generated by hand and bounding boxes were used to denote the pylons. Additionally, exclusion areas, that is, regions containing thin objects not associated with cables—flagpoles, cranes, communication towers—were marked by bounding boxes. We did not use them to create training data and excluded them from the pylon detection performance evaluation.
@@ 4.2 Baselines
PIDNet-S [84] – cable detection. This is a member of a family of models designed for semantic segmentation. We train it without boundary-aware loss components, thus also without D branch supervision. We did this because these components are mostly useful for multi-class semantic segmentation with multiple, potentially confusing, object boundaries in the image. These issues do not arise for power line cables. We start from the ImageNet-pretrained checkpoint of the PIDNet-S model and add at the end of it a single $2\times 2$ strided convolution layer to match the downsampling factor of our models. We also replaced the SGD optimizer with Adam [85] and did not use the poly learning rate strategy.
RT-DETR [27] – pylon detection. This object detector uses the ResNet18 backbone [86]. As PIDNet-S, it has been designed with real-time applications in mind. It builds on top of transformer-based object detectors. We additionally combine it with SAHI [73], training and inference strategy for object detection in high-resolution images. We evaluate two variants of patched input scales. In both cases, we run the object detector over full-resolution image and its split into $1024\times 1024$ overlapping patches, appropriately resizing the inputs for the detector and combining all of its outputs. In the second variant we additionally use similarly generated sets of $512\times 512$ and $256\times 256$ patches and add their detection results to the output. For evaluation, we rescale the final predicted bounding boxes to the downsampled output resolution of our model, use them to generate binary segmentation masks and convert them into distance masks. Also, for this model we initialize it with the pretrained weights provided with the official model implementation, obtained from training on the Objects365 [87] dataset and subsequent fine-tuning on the COCO [88] dataset.